A shiny metallic Christmas ball ornament 5.10 cm in diameter is used as a mirror. What is the mirror's focal length? please answer in cm. thank you!
1
Expert's answer
2011-06-20T12:59:59-0400
d = 5.1 cm r=d/2=2.55 cm
Shiny metallic ball is a curved mirror.
For the curved mirror F= -r/2 where r is radius of curvature;
So
F=-2.55/2=-1.275 cm.
Answer: -1.275 cm.